THE MICROWAVE ROTATIONAL SPECTRUM OF THE Ne-N_{2}O VAN DER WAALS COMPLEX

\underline{MWANIKI S. NGARI} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB, Canada, T6G 2G2.

Rotational spectra of six isotopomers of the van der Waals dimer Ne-N_{2}O were measured in the frequency range from 5 to 18 GHz using a pulsed beam cavity Fourier transform spectrometer. The spectra indicate that the complex is a prolate near symmetric rotor with a T-shaped structure. The Ne atom is on average closer to the O atom than to the terminal N atom. Both a- and b-type transitions were measured. The nuclear quadrupole hyperfine pattern due to two ^{14}N nuclei were resolved. Rotational and centrifugal distortion constants were determined, as well as the nuclear quadrupole coupling constants χ_{aa}(1), χ_{bb}(1), χ_{aa}(2), χ_{bb}(2). Effective values for the distance from the center of mass of the N_{2}O subunit to the Ne atom, and for the angle between this distance and the N_{2}O axis were obtained.